# Towards Large-Scale In-Context Reinforcement Learning by Meta-Training in Randomized Worlds

### v2

@@ 4.3 Emergence of General-Purpose ICRL by Increasing Task Number
We validate task diversity’s critical role in ICRL via independent meta-training across four datasets, each $\mathcal{D}(\mathcal{T}_{tra}(16,5))$ containing $128K$ sequences but differing in task numbers ($|\mathcal{T}_{tra}|\in\{100,1K,10K,128K\}$). Note that different trajectories can be generated from a single task, arising from the diverse behavior policies and random sampling in both decision and transition. We examine how the validation losses $\mathcal{L}_{t}$ on both seen and unseen tasks change with the number of meta-training iterations (outer-loop steps) and steps in context $t$ (inner-loop steps) simultaneously; the results are shown in Figure 7. We remark the following observations:
Task number is crucial to the generalization of ICRL. The previous investigation on ICL [26, 29] emphasizes the importance of “burstiness”. Our results demonstrate for the first time that even when using “bursty” sequences alone, both the number of tasks and their overall diversity remain critically important. Specifically, in groups with $|\mathcal{T}_{\text{train}}|\leq 10K$, over-training leads to a transiency of ICL [19] in unseen tasks but a continued improvement in seen tasks. These findings confirm and extend the discovery of the “task identification” phase mentioned in Kirsch et al. [32], Pan et al. [75]. Drawing on the theories of ICL and IWL in Chan et al. [31], a possible explanation is that IWL dominates performance, with the model memorizing tasks and ICL selecting the correct one, leading to fast seen-task adaptation but poor unseen generalization. As the number of tasks increases continuously, the model becomes more dependent on ICL since memorizing task-specific information becomes less feasible. This is characterized by the improved generalization to unseen tasks and longer adaptation periods in both seen and unseen tasks, as shown in Figure 7.
Generalization of ICRL can be at odds with its zero-shot or few-shot performance. Our results highlight a key insight on ICRL evaluation. Most previous ICRL works assess performance based on the average results over a fixed, short context span. However, our findings indicate that more generalized in-context learners may actually perform worse in zero-shot and even few-shot evaluations, particularly when there is significant overlap between the training and evaluation sets, i.e., when evaluation sets are closer to seen tasks. Therefore, we argue that it is more critical to focus on the asymptotic performance of a learner. This can be effectively evaluated by examining the performance at the final steps or episodes of a sufficiently long context, rather than short-term metrics.
@@
Validation set Metrics $|\mathcal{T}_{tra}|$
@@
Figure 7: Position-wise validation losses ($\mathcal{L}_{t}$, where lower values indicate better performance) and their properties on both seen and unseen tasks across meta-training iterations, varying context lengths, and variant number of tasks $|\mathcal{T}_{tra}|$. Each of the 4 groups of training data had $128K$ sequences, which were generated from $100,1K,10K$, and $128K$ tasks, respectively. Each dataset underwent meta-training for up to $80K$ iterations. In the table, the notation “$>$” indicates values that could not be fully determined due to training being stopped early when performance on unseen tasks began to deteriorate. The normalized gain of ICL is defined as $d_{t}=1-\mathcal{L}_{t}/\mathcal{L}_{0}$, representing the improvement of performance as the context ($t$) increases. The table summarizes key findings of this study: as the number of tasks increases, the minimum step cost required to achieve an $80\%$ normalized gain of ICL also increases.
@@ 5 Conclusions and Discussions
We propose a scalable, low-structural-bias task set for benchmarking and investigating Inverse Contextual Reinforcement Learning (ICRL). Our framework introduces two key innovations: stepwise supervision and prior information induction, creating a more efficient ICRL system. Experimental results demonstrate that our model achieves superior generalization across RL tasks compared to existing approaches. The proposed model generalizes to a broader range of RL tasks than ever before. Our work introduces a more scalable testbed for ICRL research.
Broader impact: Complementing prior studies, our findings highlight that task diversity and sequence length—alongside sequence burstiness—are key determinants of general ICRL. Our results also advocate shifting evaluation metrics toward asymptotic performance measures. This work further motivates the construction of carefully curated synthetic datasets specifically designed for large-scale meta-training.
Limitations and future work: We acknowledge that the discrete state and action space is the most significant limitation of this work, as it restricts its application to environments with continuous state and action spaces. Extending this work to continuous and partially observable environments, as well as incorporating additional modalities, would significantly broaden its scope of application.
@@ Appendix A Notations
SS Step-wise Supervision
@@
Appendix B Details of AnyMDP Environments
B.1 Procedural generation of high-quality MDPs
Algorithm 1 AnyMDP $TaskSampler$
1:  Input: $n_{s}$, $n_{a}$ Returns: $\tau(n_{s},n_{a})$
2:  randomly generate: an ordered list arranged from low-valued to high-valued states $S=[s_{1},s_{2},...,s_{n_{s}}]$ by permute $\mathcal{S}=\{1,...,n_{s}\}$
3:  randomly sample: states for reset $\mathcal{S}_{0}\subset\{s_{0},s_{1},...,s_{b_{0}}\}$
4:  randomly sample: states triggering termination $\mathcal{S}_{E}\subset\mathcal{S}/\mathcal{S}_{0}$
5:  for $i\leftarrow 1$ to $n_{s}$ do
6:     randomly sample: $\hat{\mathcal{P}_{\mathfrak{r}}}(\cdot|s_{i})$, s.t.
7:                                    $\forall j>i+b_{+}$ and $\forall j<i-b_{-}$, $\hat{\mathcal{P}_{\mathfrak{r}}}(s_{j}|s_{i})=0$ and
8:                                    $\sum_{j<i}\hat{\mathcal{P}_{\mathfrak{r}}}(s_{j}|s_{i})>\eta$ if $i>b_{-}$ and
9:                                    $\sum_{j\leq i}\hat{\mathcal{P}}_{\mathfrak{r}}(s_{i+1}|s_{j})>\epsilon$ if $i<n_{s}-1$
10:     randomly sample: $w(s_{i},a_{k})$ $\forall k\in[1,n_{a}]$
11:     set: $\mathcal{P}(s_{j}|s_{i},a_{k})=\hat{\mathcal{P}_{\mathfrak{r}}}(s_{j}|s_{i})\exp{w(s,a_{k})}/\sum_{a}\exp(w(s,a))$ $\forall j,k$
12:  end for
13:  randomly sample: state-action dependent rewards $r_{sa}\in\mathbb{R}^{n_{s}\times n_{a}}$
14:  randomly sample: potential $v_{s}\in\mathbb{R}^{n_{s}}$
15:  repeat
16:     randomly sample: state-wise reward function $r_{s}\in\mathbb{R}^{n_{s}}$, s.t. $\forall j>i,r_{s}(s_{j})\geq r_{s}(s_{i})$
17:     set:$\mu_{\mathcal{R}}(s,a,s^{\prime})=r_{sa}(s,a)+r_{s}(s^{\prime})+v_{s}(s)-v_{s}(s^{\prime})$
18:     calculate: value function $V^{*}(s)$ based on $\mathcal{P},\mathcal{R},\mathcal{S}_{0},\mathcal{S}_{E}$
19:  until $V^{*}(s_{g})>\max_{s\in\mathcal{S}_{0}}V^{*}(s)+b_{V}$
20:  final validate: calculate oracle policy steady state distribution ($p_{\mathfrak{o}}$)
21:                if $-\sum_{s}p_{\mathfrak{o}}\log{p_{\mathfrak{o}}}/\log{n_{s}}>H_{0}$ then accept, else resample the task.
To effectively sample the tuple $\langle\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},\mathcal{S}_{0},\mathcal{S}_{E},T_{E},\gamma\rangle$ and ensure broad coverage of the task space while providing sufficient challenges for resolution, we propose a procedural generation pipeline, which is detailed in Algorithm 1. This approach is guided by the general principle that states which are more difficult to reach tend to offer greater rewards. The procedural generation methods emphasize two key features: a banded transition kernel and a composite reward function.
Banded transition kernel: We first demonstrate that this sampling method ensures the reachability of all states while simultaneously guaranteeing that the probability of finding a solution through random exploration decreases at least exponentially.
We focus on the Markov chain (MC) defined by running a uniform random policy on the MDP, where actions are chosen by: $a\sim\pi^{\mathfrak{r}}(a|s)=\frac{1}{|\mathcal{A}|}$. This Markov chain can be formally represented as $\langle\hat{\mathcal{P}}^{\mathfrak{r}},\mathcal{S}_{0},\mathcal{S}_{E}\rangle$, with $\hat{\mathcal{P}}_{\mathfrak{r}}$ representing the state-state transition kernel as follows:
$\hat{\mathcal{P}}_{\mathfrak{r}}(s^{\prime}|s)=\mathbb{E}_{a\sim\pi^{\mathfrak{r}}(a|s)|}\mathcal{P}(s^{\prime}|s,a)=P_{\mathfrak{r},s,s^{\prime}}$ (6)
We further define the stationary distribution (SD) of the Markov chain $P_{\mathfrak{r}}$ as follows:
$p_{\mathfrak{r}}P_{\mathfrak{r}}=p_{\mathfrak{r}}$ (7)
Here $p_{\mathfrak{r}}\in\mathbb{R}^{n_{s}}$ is a probability vector that sums to 1, representing the asymptotic distribution over the state space under the transition matrix $P_{\mathfrak{r}}$. With this foundation, we can prove the following:
Theorem 1.
For $j>b$, where $b=\max(b_{-},b_{0})+b_{+}$, there exists a value $0<\delta<1/(b_{+}+1)$. If $\eta>1-\delta$,the sampling algorithm described in Algorithm 1 ensures that the random policy Markov chain has a stationary distribution (SD). Specifically, for the state $s_{j}$, the SD $p_{\mathfrak{r}}(s_{j})$ satisfies the following bounds:
$C_{1}\epsilon^{j-b}<p_{\mathfrak{r}}(s_{j})<C_{2}(1-\delta)^{j-b}$,
where $C_{1}$ and $C_{2}$ are positive constants.
Proof.
@@ B.1 Procedural generation of high-quality MDPs
First, we assume that $p_{+,i,i}/p_{+,i,i-1}=1-\delta$, Equation 8. From this assumption, we derive the following:
@@
By applying the conditions $0<\delta<1/(b_{+}+1)$ and $\eta>1.0-\delta$, we can further bound the expression in Equation 9 as follows:
@@
Equation 10 proves that $p_{+,s_{i},s_{j}}$ decays at a rate faster than $(1-\delta)$ as $j$ increases. Consequently, the SD $p_{\mathfrak{r}}$ also decays faster than $(1-\delta)$.
@@
By analyzing this construction, we can validate that $p_{-,s_{i},s_{j}}$ decays at a rate slower than $\epsilon$ when $j$ increases. Consequently, the SD $p_{\mathfrak{r}}$ also decays slower than $\epsilon$. This completes the proof of Theorem 1. ∎
Theorem 1 guarantees two critical properties. Firstly, the transition kernel remains irreducible, preventing isolated states. Secondly, it ensures that the states are increasingly difficult to reach as $j$ increases, akin to the concept of “higher-hanging fruits.”.
While this theoretical bound provides valuable non-triviality guarantees, it remains inherently conservative due to the analytical challenges posed by randomly generated task structures. Empirically, we observe that setting $b_{+}\leq n_{s}/4$, $b_{-}\geq n_{s}/2$, $\epsilon>1.0e-3$, $\eta>0.5$ is enough to consistently yield high-quality Markov chain formulations.
Composite reward function: The reward function is designed to ensure that the value function increases progressively from low-valued states to high-valued states (in the direction of increasing $j$ in $S=[s_{1},s_{2},...,s_{n_{s}}]$). Specifically, $\mathcal{R}(s,a,s^{\prime})$ is composed of the following three components:
@@
  • Randomly sampled state-action dependent reward function $r_{sa}(s,a)$.
  • Randomly sampled potential-based reward function $v_{s}(s)-v_{s}(s^{\prime})$. This can be interpreted as a form of random reward shaping.
  • State-wise reward function $r_{s}(s^{\prime})$, which is monotonically increasing with $j$. This is the most crucial reward function for distinguishing high-value states from low-value states.
To ensure non-trivial task formulations, we implement a suite of validation checks. A critical safeguard involves enforcing a minimum threshold on the normalized entropy $\mathcal{H}(p_{\mathfrak{o}})$ of the SD under the oracle policy $\mathfrak{o}$. For the oracle policy that might end up in absorbing states, this constraint prevents degenerate solutions to a single high-value state.
B.2 Supplementary Empirical Study on AnyMDP
@@
Figure 8 illustrates the computational cost of generating AnyMDP tasks for various state space sizes $n_{s}\in\{8,16,32,64,128\}$. Notably, $n_{s}=8$ exhibits significantly higher computation times. This is primarily due to the frequent resampling required when the value function check fails. It is important to note that the AnyMDP task generation process was executed on single CPU. Given this, the use of readily available parallelization techniques could significantly accelerate task generation.
@@
Figure 9 illustrates the performance of Tabular Q-Learning and Proximal Policy Optimization (PPO) on AnyMDP tasks with varying state space and action space sizes. The results indicate that increasing either the state space size ($n_{s}$) or the action space size ($n_{a}$) enhances the complexity of the task, as evidenced by the need for more training steps to achieve convergence. OmniRL’s result on AnyMDP tasks with different state spaces also supports this phenomenon, shown in Figure 10. Additionally, an increase in the state space size ($n_{s}$) leads to a higher number of steps per episode.
Figure 11: An ablation study comparing AnyMDP tasks with tasks sampled using other methods demonstrates that the procedural generation algorithm (Algorithm 1) produces tasks of higher quality.
Figure 11 compares three types of tasks: tasks sampled from Algorithm 1, tasks where $\langle\mathcal{P},\mathcal{R}\rangle$ are directly sampled from naive Gaussian distributions, and AnyMDP tasks without composite reward (where $\mathcal{R}$ is sampled from a Gaussian distribution). The results demonstrate that AnyMDP tasks provide greater challenges for reinforcement learning mechanisms.
B.3 Visualizing Discrete MDPs
@@
(d) FrozenLake(non-slippery)
(e) FrozenLake(Slippery))
(f) Discrete-Pendulum (g=9.8)
(g) DarkRoom($8\times 8,[7,7]$)
(h) DarkRoom($10\times 10,[9,9]$)
@@
Figure 12: This figure visualizes three tasks sampled from AnyMDP, with the number of states $n_{s}$ varying across $\{16,32,64\}$, Gymnasium tasks: non-slippery and slippery FrozenLake and Discrete-Pendulum, and 3 tasks sampled from DarkRoom with different dimension and goals. States are reordered according to the SD of $p_{\mathfrak{r}}$, ordered from high to low. Gray blocks indicate state transition kernels. Red and blue blocks mark pitfalls ($\mathcal{S}^{-}_{E}$) and goals ($\mathcal{S}^{+}_{E}$), respectively, which trigger episode termination. Green blocks mark $\mathcal{S}_{0}$. The black line denotes the state value function under the optimal policy $V^{*}(s)$. Notably, AnyMDP is capable of generating a diverse range of tasks, including those with and without pitfalls and goals. The visualizations demonstrate that tasks generated by AnyMDP can be of comparable quality to those designed by humans. A common principle observed is that higher rewards are often associated with more challenging goals, akin to the concept of “high-hanging fruit”.
Following the previous analysis, for any discrete Markov Decision Process (MDP), we can rearrange the states such that the SD $p_{\mathfrak{r}}$ decreases monotonically. We then plot the transition kernel $P_{\mathfrak{r}}$ in this rearranged order. In the visualization, we use varying opacity to represent the elements of $P_{\mathfrak{r}}$ and different colors to distinguish the initial states $\mathcal{S}_{0}$, positively rewarded terminal states (goals) $\mathcal{S}^{+}_{E}$, and negatively rewarded terminal states $\mathcal{S}^{-}_{E}$. This visualization, shown in Figure 12, enables us to analyze both procedurally generated AnyMDP tasks and human-designed Gymnasium tasks. Several interesting observations can be made:
@@
  • Banded transition kernel. When ordered by decreasing SD probability, the transition kernels of all Markov chains display the characteristics of a banded matrix. This observation further validates the effectiveness of the procedural generation method outlined in Algorithm 1.
Appendix C Details of Experiment Settings

### v3

@@ 5.3 Emergence of General-Purpose ICRL by Increasing Task Number
We validate task diversity’s critical role in ICRL via independent meta-training across four datasets, each $\mathcal{D}(\mathcal{T}_{tra}(16,5))$ containing $128K$ sequences but differing in task numbers ($|\mathcal{T}_{tra}|\in\{100,1K,10K,128K\}$). Note that different trajectories can be generated from a single task, arising from the diverse behavior policies and random sampling in both decision and transition. We examine how the validation losses $\mathcal{L}_{t}$ on both seen and unseen tasks change with the number of meta-training iterations (outer-loop steps) and steps in context $t$ (inner-loop steps) simultaneously; the results are shown in Figure˜7. We remark the following observations:
Task scale and diversity are crucial to the generalization of ICRL. The previous investigation on ICL [25, 28] emphasizes the importance of “burstiness”. Our results demonstrate for the first time that even when using “bursty” sequences alone, both the number of tasks and their overall diversity remain critically important. Specifically, in groups with $|\mathcal{T}_{\text{train}}|\leq 10K$, over-training leads to a transiency of ICL [74] in unseen tasks but a continued improvement in seen tasks. These findings confirm and extend the discovery of the “task identification” phase mentioned in Kirsch et al. [34], Pan et al. [75]. Drawing on the theories of ICL and IWL in Chan et al. [76], a possible explanation is that IWL dominates performance, with the model memorizing tasks and ICL selecting the correct one, leading to fast seen-task adaptation but poor unseen generalization. As the number of tasks increases continuously, the model becomes more dependent on ICL since memorizing task-specific information becomes less feasible. This is characterized by the improved generalization to unseen tasks and longer adaptation periods in both seen and unseen tasks, as shown in Figure˜7.
Long-context adaptation is a tax to pay for generalization scope. Our results highlight a key insight on ICRL evaluation. Most previous ICRL works assess performance based on the average results over a fixed, short context span. However, our findings indicate that more generalized in-context learners may actually perform worse in zero-shot and even few-shot evaluations, particularly when there is significant overlap between the training and evaluation sets, i.e., when evaluation sets are closer to seen tasks. Therefore, we argue that it is more critical to focus on the asymptotic performance of a learner. This can be effectively evaluated by examining the performance at the final steps or episodes of a sufficiently long context, rather than short-term metrics.
@@
Validation
set
Metrics $|\mathcal{T}_{tra}|$
@@
Figure 7: Position-wise validation losses ($\mathcal{L}_{t}$, where lower values indicate better performance) and their properties on both seen and unseen tasks across meta-training iterations, varying context lengths, and variant number of tasks $|\mathcal{T}_{tra}|$. Each of the 4 groups of training data had $128K$ sequences, which were generated from $100,1K,10K$, and $128K$ tasks, respectively. Each dataset underwent meta-training for up to $80K$ iterations. In the table, the notation “>>” indicates values that could not be fully determined due to training being stopped early when performance on unseen tasks began to deteriorate. The normalized gain of ICL is defined as $d_{t}=1-\mathcal{L}_{t}/\mathcal{L}_{0}$, representing the improvement of performance as the context ($t$) increases. The table summarizes key findings of this study: as the number of tasks increases, the minimum step cost required to achieve an $80\%$ normalized gain of ICL also increases.
@@ 6 Conclusions and Discussions
We introduce AnyMDP, a scalable, low-bias task suite for benchmarking In-Context Reinforcement Learning (ICRL), and a companion framework featuring two innovations: Decoupled Policy Distillation and prior-information induction. Trained solely on AnyMDP, our model OmniRL outperforms prior methods in generalization, mastering wider task families and supporting more diverse learning paradigms.
Broader impact: Complementing prior studies, our findings highlight that task diversity and long sequence length are key determinants of general-purpose ICRL. Our results also indicate that long context is a necessary cost for generalization, thus advocate shifting evaluation metrics toward asymptotic performance measures. This work further motivates the construction of carefully curated synthetic datasets specifically designed for large-scale meta-training.
Limitations and future work: While procedural Discrete MDPs provide clean benchmarks for ICRL and long-context modeling, their extension to continuous state–action spaces remains bottlenecked. We recognize that a central challenge in this direction is to balance task complexity, task diversity, and data integrity while simultaneously calibrating generalization scope against practical constraints.
Acknowledgments and Disclosure of Funding
This project is supported by Longgang District Shenzhen’s “Ten Action Plan” for supporting innovation projects (under Grant LGKCSDPT2024002) and National Natural Science Foundation of China (62033012).
@@ Appendix A Notations
$S$ Step-wise Supervision
@@
Appendix B Additional Information on AnyMDP
B.1 Proof of Theorem˜1
@@
First, we assume that $p_{+,i,i}/p_{+,i,i-1}=1-\delta$, Equation˜6. From this assumption, we derive the following:
@@
By applying the conditions $0<\delta<1/(b_{+}+1)$ and $\eta>1.0-\delta$, we can further bound the expression in Equation˜7 as follows:
@@
Equation˜8 proves that $p_{+,s_{i},s_{j}}$ decays at a rate faster than $(1-\delta)$ as $j$ increases. Consequently, the SD $p_{\mathfrak{r}}$ also decays faster than $(1-\delta)$.
@@
By analyzing this construction, we can validate that $p_{-,s_{i},s_{j}}$ decays at a rate slower than $\epsilon$ when $j$ increases. Consequently, the SD $p_{\mathfrak{r}}$ also decays slower than $\epsilon$. This completes the proof of Theorem˜1.
B.2 Details of Procedural Generation
Algorithm 1 AnyMDP TaskSamplerTaskSampler
1:Input: $n_{s}$, $n_{a}$ Returns: $\tau(n_{s},n_{a})$
2: Randomly generate an ordered list arranged from low-valued to high-valued states $S=[s_{1},s_{2},...,s_{n_{s}}]$ by permute $\mathcal{S}=\{1,...,n_{s}\}$
3: Randomly sample states for reset $\mathcal{S}_{0}\subset\{s_{0},s_{1},...,s_{b_{0}}\}$
4: Randomly sample states triggering termination $\mathcal{S}_{E}\subset\mathcal{S}/\mathcal{S}_{0}$ $\triangleright$ Sample banded state transition matrix under uniform random policy
5:for $i\leftarrow 1$ to $n_{s}$ do
6:  Randomly sample $\hat{\mathcal{P}_{\mathfrak{r}}}(\cdot|s_{i})$, s.t. banded transition constraints (Equation˜1)
[…]
8: Calculate $\bar{\mathcal{P}}_{\mathfrak{r}}=(\hat{\mathcal{P}}_{\mathfrak{r}})^{K}$ by considering $\mathcal{S}_{0},\mathcal{S}_{E}$; if $\mathbb{E}_{j}[\mathbb{V}_{i}(\bar{\mathcal{P}_{\mathfrak{r}}}[s_{j}|s_{i}])]>\varepsilon$, go back to resample $\hat{\mathcal{P}}_{\mathfrak{r}}$. $\triangleright$ Sample state-action transition with the constraint of the state transition
9:for $i\leftarrow 1$ to $n_{s}$ do
10:  $\forall k\in[1,n_{a}]$, randomly sample scalars $w_{i}(a_{k})\in[i-b_{-},i+{b_{+}}]$, $\sigma_{k}\sim Exp(1)$
11:  Sample state-action transition $\mathcal{P}(s_{j}|s_{i},a_{k})$ by applying Equation˜9
[…]
13: Randomly sample state-action-dependent rewards $r_{sa}\in\mathbb{R}^{n_{s}\times n_{a}}$
14: Randomly sample state-dependent potential $v_{s}\in\mathbb{R}^{n_{s}}$
15:repeat
@@ B.2 Details of Procedural Generation
16:  Randomly sample state-dependent reward function $r_{s}\in\mathbb{R}^{n_{s}}$, s.t. $\forall j>i,r_{s}(s_{j})\geq r_{s}(s_{i})$
17:  Set $\mu_{\mathcal{R}}(s,a,s^{\prime})$ by composite rewards (Equation˜10)
18:  Calculate value function $V^{*}(s)$ based on $\mathcal{P},\mathcal{R},\mathcal{S}_{0},\mathcal{S}_{E}$
19:until Ascending value function condition is satisfied (Equation˜2) $\triangleright$ Ensure the optimal trajectory is not trivial
20:Final validation: calculate oracle policy steady state distribution ($p_{\mathfrak{o}}$)
21:       if $\sum_{s}p_{\mathfrak{o}}\log{p_{\mathfrak{o}}}/\log{n_{s}}>H_{0}$ then accept, else resample the task.
Algorithm˜1 elaborates on the detailed procedural generation of AnyMDP tasks.
Transition Sampling. The generation process commences by randomly sampling a ranking of states, denoted as $s_{1},...,s_{n_{s}}$. The initial state set $\mathcal{S}_{0}$ are sampled from $\{s_{j}|j<n_{0}\}$, while the absorbing state set (including pitfalls and goals) $\mathcal{S}_{E}$ is randomly sampled. The goals are constrained only to $s_{n_{s}}$. Subsequently, the average transition matrix is sampled, and based on this matrix, it is randomly decomposed into a state-action transition matrix. Notice that the state-action transition is constructed utilizing the average state transition kernel and the weights $w_{i}(a_{k})$, as follows:
$\mathcal{P}(s_{j}|s_{i},a_{k})=\hat{\mathcal{P}_{\mathfrak{r}}}(s_{j}|s_{i})\frac{\exp{[-(w_{i}(a_{k})-j)^{2}/\sigma_{k}^{2}}]}{\sum_{l}\exp{[-(w_{i}(a_{l})-j)^{2}/\sigma_{l}^{2}}]}$ (9)
Ergodicity in the average transition of AnyMDP is checked by computing the stationary distribution for all initial states and evaluating the variance among them. In practice, we observed that the majority of sampled transition matrices adhering to banded transition constraints exhibit ergodicity, with non-ergodic scenarios being exceedingly uncommon.
While the theoretical bound in Theorem˜1 provides valuable non-triviality guarantees, it remains inherently conservative due to the analytical challenges posed by randomly generated task structures. Empirically, we observe that setting $b_{+}\leq n_{s}/4$, $b_{-}\geq n_{s}/2$, $\epsilon>1.0e-3$, $\eta>0.5$ is enough to consistently yield high-quality Markov chain formulations.
Composite Reward Sampling. To further enhance the quality of the task, rather than naively sampling the reward function $\mathcal{R}$ from $\mathbb{R}^{n_{s}\times n_{a}\times n_{s}}$, e independently sample the mean reward matrix $\bf{\mu}^{R}$ and the noises $\Sigma^{R}$. This ensures that the reward $r(s,a,s^{\prime})\sim\mathcal{N}(\mu^{R}_{s,a,s^{\prime}},\Sigma^{R}_{s,a,s^{\prime}})$. The mean reward matrix $\mu_{R,s,a,s^{\prime}}$ is sampled from composite reward functions, which is defined as follows:
$\mu^{R}_{s,a,s^{\prime}}=r^{s}(s^{\prime})+r^{sa}(s,a)+v^{s}(s)-v^{s}(s^{\prime})$ (10)
We first sample the state-dependent reward $r^{s}$ based on the ranking of states $s_{1},...,s_{n_{s}}$ and whether each state corresponds to pitfalls or goals. Subsequently, we sample the disturbance reward function $r^{sa}$ and state-value function $v^{s}$, which can be interpreted as random state-action costs and random reward shaping terms, respectively.
To ensure non-trivial task formulations, we implement a suite of validation checks, including: 1. Ergodicity check on the average transition function; 2. Ascending value function check to guarantee monotonic improvement; 3. Minimum threshold check on the normalized entropy $\mathcal{H}(p_{\mathfrak{o}})$ of the state distribution (SD) under the oracle policy $\mathfrak{o}$. If any of these checks fail, we either resample the transition dynamics or readjust the sampling of composite rewards. Details of the sampling process are provided in the Appendices.
B.3 Additional Properties of AnyMDP
@@
Computational Efficiency of Task Sampling. Figure˜8 illustrates the computational cost of generating AnyMDP tasks for various state space sizes $n_{s}\in\{8,16,32,64,128\}$. Notably, $n_{s}=8$ exhibits significantly higher computation times. This is primarily due to the frequent resampling required when final validation check fails. It is important to note that the AnyMDP task generation process was executed on single CPU. Given this, the use of readily available parallelization techniques could significantly accelerate task generation.
@@
Relations between solution difficulty and the state–action space. Figure 9 illustrates the performance of Tabular Q-Learning and Proximal Policy Optimization (PPO) on AnyMDP tasks with varying state space and action space sizes. The results indicate that increasing either the state space size ($n_{s}$) or the action space size ($n_{a}$) enhances the complexity of the task, as evidenced by the need for more training steps to achieve convergence. OmniRL’s result on AnyMDP tasks with different state spaces also supports this phenomenon, shown in Figure 10. Additionally, an increase in the state space size ($n_{s}$) leads to a higher number of steps per episode.
@@
(d) Garnet(64,5,2)
(e) FrozenLake(non-slippery)
(f) FrozenLake(Slippery))
(g) Discrete-Pendulum (g=9.8)
(h) DarkRoom($8\times 8,[7,7]$)
@@
Figure 11: Visualization of three tasks sampled from AnyMDP, with the number of states $n_{s}$ varying across $\{16,32,64\}$, Garnet(64,5,2)(64,5,2), and Gymnasium tasks including non-slippery and slippery FrozenLake and Discrete-Pendulum, and 2 tasks sampled from DarkRoom. States are reordered according to the SD ($p_{\mathfrak{r}}$), ordered from high to low. Gray blocks indicate positive state transition kernels. Red and blue blocks mark pitfalls ($\mathcal{S}^{-}_{E}$) and goals ($\mathcal{S}^{+}_{E}$), respectively, which trigger episode termination. Green blocks mark $\mathcal{S}_{0}$. The black line denotes the state value function under the optimal policy $V^{*}(s)$. Notably, AnyMDP is capable of generating a diverse range of tasks, including those with and without pitfalls and goals. The visualizations demonstrate that tasks generated by AnyMDP can be of comparable quality to those designed by humans. A common principle observed is that higher rewards are often associated with more challenging goals, akin to the concept of “high-hanging fruit”.
Visualizing discrete MDPs. Following the previous analysis, for any discrete Markov Decision Process (MDP), we can rearrange the states such that the SD $p_{\mathfrak{r}}$ decreases monotonically. We then plot the transition kernel $P_{\mathfrak{r}}$ in this rearranged order. In the visualization, we use varying opacity to represent the elements of $P_{\mathfrak{r}}$ and different colors to distinguish the initial states $\mathcal{S}_{0}$, positively rewarded terminal states (goals) $\mathcal{S}^{+}_{E}$, and negatively rewarded terminal states $\mathcal{S}^{-}_{E}$. This visualization, shown in Figure˜11, enables us to analyze both procedurally generated AnyMDP tasks and human-designed Gymnasium tasks. Several interesting observations can be made:
@@
  • Banded transition kernel. When ordered by decreasing SD probability, the transition kernels of all Markov chains display the characteristics of a banded matrix. This observation further validates the effectiveness of the procedural generation method outlined in Algorithm˜1.
Appendix C Additional Information on Experiment Settings